OpenStudy (anonymous):

Please help me solve this problem Rationalize the denominator to the problem 2 to the square root of v divided by 2-square root of z

OpenStudy (anonymous):

are u asking about this : \[\frac{2^{\sqrt{v}}}{2-\sqrt{z}}\]

OpenStudy (anonymous):

multiply by 2+sqrt(z) to numerator and denominator

OpenStudy (anonymous):

I messed up. They are both z.

OpenStudy (anonymous):

\[\frac{2^{\sqrt{z}}\times (2+\sqrt{z})}{(2-\sqrt{z})\times(2+\sqrt{z})}\] \[=\frac{2^{\sqrt{z}}\times(2+\sqrt{z})}{4-z}\]